The bill for 5 glasses of apple juice and 4 salads is $9.50,but the bill for 4 glasses of apple juice and 5 salads is $ 10.30. What would be the bill for a glass of juice and a salad?
Found 2 solutions by mananth, lwsshak3:
Answer by mananth(16949)

Apple juice cost x
Salad cost ------------ y
5x+4y=9.5 .............1
4x+5y=10.3 .............2
Eliminate y
multiply (1)by -5
Multiply (2) by 4
-25x-20 y=-47.5
16x+20y =41.2
Add the two equations
-9x=-6.3
/-9
x=0.70
plug value of x in (1)
5x+4y=9.5
3.5+4y=9.5
4y=9.5-3.5
4y=6
y=1.50

Answer by lwsshak3(11628)

The bill for 5 glasses of apple juice and 4 salads is $9.50,but the bill for 4 glasses of apple juice and 5 salads is $ 10.30. What would be the bill for a glass of juice and a salad?
..
let x=cost of one glass of apple juice
let y=cost of one salad
5x+4y=9.50
4x+5y=10.30
..
20x+16y=38.00
20x+25y=51.50
..
-9y=-13.50
y=13.50/9=1.50
4x=10.30-5y=10.30-5*1.5=10.30-7.50=2.80
4x=2.80
x=.70
x+y=1.50+.70=2.20
Ans:
The bill for a glass of juice and a salad=$2.20